OpenStudy (anonymous):

“The captain of a rafting boat takes his crew down river for 6 hours, covering a total of 56 miles. When he realizes he left one of his passengers behind, he is forced to paddle back upriver for 8 miles. The trip back up took him 3 hours.” A student used the following system of equations to model the situation given above. Which of the statements is true about the system of equations the student used? b = the speed of the boat r = the speed of the current {8(b-r)=3 {6(b+r)=56 A.The system of equations the student used is correct. B.The top equation in the system is incorrect.

OpenStudy (anonymous):

C. The bottom equation in the system is incorrect. D. Both equations in the system are incorrect.

OpenStudy (kendricklamar2014):

http://openstudy.com/study#/updates/52b39d4ae4b02932bbbf0331

OpenStudy (agent0smith):

{8(b-r)=3 this should be 3(b-r) = 8 since the distance is 8 miles, the speed is b-r, and time is 3 hours. d = rt, or distance = rate multiplied by time

OpenStudy (anonymous):

so the bottom equation is incorrect. OR the equations the student used or both correct???

OpenStudy (kendricklamar2014):

the bottom equation is incorrect

OpenStudy (agent0smith):

Read what I wrote. The TOP equation is the one i corrected.